OpenStudy (anonymous):

Given the function g(x) = 4x − 5, find g(−3).

OpenStudy (saifoo.khan):

g(x) = 4x − 5, g(-3) = 4(-3) − 5

OpenStudy (saifoo.khan):

-12 - 5 -17

jimthompson5910 (jim_thompson5910):

g(x)=4x-5 g(-3)=4(-3)-5 g(-3)=-12-5 g(-3)=-17

OpenStudy (anonymous):

g (-3) = 4 (-3) -5 = -17